Calculate the acceleration of gravity relative to earth's acceleration (i.e., 1 g) at the surface of Venus. The mass of Venus is 4.87 x 10²⁴ kg and the planet's radius is 6.05 x 10⁶ m. The universal gravitational constant = G = 6.67 x 10⁻¹¹ N*m²/kg².


1
Expert's answer
2021-03-29T09:01:51-0400
"g_V=\dfrac{GM_M}{R_M^2},""g_V=\dfrac{6.67\cdot10^{-11}\ \dfrac{Nm^2}{kg^2}\cdot4.87\cdot10^{24}\ kg}{(6.05\cdot10^6\ m)^2},""g_V=8.87\ \dfrac{m}{s^2}\cdot\dfrac{1\ g}{9.8\ \dfrac{m}{s^2}}=0.9g."